Professor Wendy Smith has been offered the following​ opportunity: A law firm would like to retain her for an upfront payment of

$50,000.

In​ return, for the next year the firm would have access to eight hours of her time every month. As an alternative payment​ arrangement, the firm would pay Professor​ Smith's hourly rate for the eight hours each month. ​ Smith's rate is

$550

per hour and her opportunity cost of capital is

15%

per year. What does the IRR rule advise regarding the payment​ arrangement? (Hint: Find the monthly rate that will yield an effective annual rate of

15%​.)

What about the NPV​ rule?
Answer #1

IRR can be calculated using I/Y function on a calculator

N = 12, PMT = 550 x 8 = 4,400, PV = -50,000, FV = 0

=> Compute I/Y = 0.85% (monthly)

IRR = 0.85% x 12 = 10.18%

As the IRR of monthly option is less, she should opt for upfront payment.

NPV can be calculated using PV function

N = 12, I/Y = 15%/12, PMT = 4400, FV = 0

=> Compute PV = $48,748.97

NPV = 48,748.97 - 55,000 = -1,251.03

As NPV is negative, monthly option should not be opted.